# Burnside type results for fusion categories

### v10

@@ Abstract.
In this paper, we extend a classical vanishing result of Burnside from the character tables of finite groups to the character tables of commutative fusion rings, or more generally to a certain class of abelian normalizable hypergroups. We also treat the dual vanishing result. We show that any nilpotent fusion categories satisfy both Burnside’s property and its dual. Using Drinfeld’s map, we obtain that the Grothendieck ring of any weakly-integral modular fusion category satisfies both properties. As applications, we prove new identities that hold in the Grothendieck ring of any weakly-integral fusion category satisfying the dual-Burnside’s property, thus providing new categorification criteria. In particular we improve [OY23, Theorem 4.5] as follows: A weakly integral modular fusion category of $\mathrm{FPdim}$ $md$ with $d$ square-free coprime with $m$ and $\mathrm{FPdim}(X)^{2}$ for every simple object $X$, has a pointed modular fusion subcategory of $\mathrm{FPdim}$ $d$. We also prove some new results on the perfect fusion categories.
@@
18M20; 16T30
The first author is supported by a grant of the Ministry of Research, Innovation and Digitization, CNCS/CCCDI - UEFISCDI, project number PN-III-P4-ID-PCE-2020-0878, within PNCDI III. The second author is supported by BIMSA Start-up Research Fund and Foreign Youth Talent Program from the Ministry of Sciences and Technology of China.
@@ 1. Introduction
A classical result of Burnside in the character theory of finite groups states that any irreducible non-linear character of a finite group vanishes on at least one element of the group. This can be stated as follows: in the character table of a finite group, the row of every character of degree $\neq 1$ contains a zero entry.
More recently, the dual version of this result was also intensively studied in the literature. This dual version describes the group elements that vanish on at least one irreducible character, see [INW99] and the references therein. Burnside’s result was extended to all weakly integral fusion categories, initially in the context of modular categories in [GNN09, Appendix], and subsequently in the general framework in [Bur23].
@@
Recall that in [Bur21] the author introduced the notion of dual of a Grothendieck ring. This was achieved based on work of Harrison [Har79] on dualizable probability groups and on [ZZ21]. More generally, this notion of dual of a fusion ring was extended to arbitrary fusion rings in [BP22]. It was noticed in [Bur21] that the dual of the Grothendieck ring of a pivotal fusion category is isomorphic to the center of the category as defined in [Shi17].
@@
(1.1) $aa^{*}=N_{a,a^{*}}^{1}1\text{ (in other words, }N_{a,a^{*}}^{c}=\delta_{1,c}N_{a,a^{*}}^{1}\text{)}$
is called a grouplike element of ${(H,\mathcal{B})}$.
@@ 1.1. The dual hypergroup of a normalized hypergroup
Let ${(H,\mathcal{B})}$ be an abelian normalized hypergroup. By the abelian assumption, $^{\widehat{\mathcal{B}}}$ forms a basis for the dual $H^{*}$, which then is $\mathbb{C}[{\widehat{\mathcal{B}}}]$, and on which one can define a multiplication. On the basis $^{\widehat{\mathcal{B}}}$, the multiplication $\mu_{i}\star\mu_{j}$ is defined by declaring
@@
and then extending linearly $\mu_{i}\star\mu_{j}$ on the entire ${\mathbb{C}[\mathcal{B}]}$. In this way, one obtains an algebra structure on $H^{*}$.
@@
By [Bur21, Proposition 2.3], for any abelian normalized hypergroup $(H,{\mathcal{B}})$, the pair $(H^{*},{\widehat{\mathcal{B}}})$ is also an abelian normalized hypergroup. Moreover, the involution on $^{\widehat{\mathcal{B}}}$ is given by $\mu_{j}\mapsto\mu_{j^{\#}}$, where $\mu_{j^{\#}}\in{\widehat{\mathcal{B}}}$ is defined by $\mu_{j^{\#}}(a):=\mu_{j}(a^{*})$ for all $a\in\mathcal{B}$. We denote by $\widehat{{(H,\mathcal{B})}}:=(H^{*},{\widehat{\mathcal{B}}})$ the dual hypergroup of the abelian normalized hypergroup ${(H,\mathcal{B})}$. It is also well known that for a normalized hypergroup ${(H,\mathcal{B})}$, then $\mu_{j}(a^{*})=\overline{\mu_{j}(a)}$ for all $a\in\mathcal{B}$. We also use the notation ${\widehat{h}}_{j}:={\widehat{p}}_{1}(j,j^{\#})^{-1}$, called the order of $\mu_{j}$ in the dual hypergroup ${\widehat{(H,\mathcal{B})}}$, see Lemma 2.6.
For any RN-hypergroup ${(H,\mathcal{B})}$, we denote by ${(H,\mathcal{B})}_{ad}$ its adjoint sub-hypergroup, see §6.2. As for fusion rings, this is defined as the sub-hypergroup of ${(H,\mathcal{B})}$ generated by the element $I(1):=\sum_{i=1}^{m}h_{i}x_{i}x_{i^{*}}$, where $h_{i}:=1/N^{1}_{i,i^{*}}$.
@@
A hypergroup ${(H,\mathcal{B})}$ is called normalizable if one obtains a normalized hypergroup after rescaling it. (see §2.2 for more details.)
@@ 1.2. Main results of the paper
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup such that ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN. The following are equivalent (see Definition 2.4 for $n(\_)$):
  1. (1)
    The hypergroup ${(H,\mathcal{B})}$ is Burnside.
  2. (2)
    The following identity holds on ${\widehat{(H,\mathcal{B})}}$:
@@
Note that the right hand side of the above equation coincides to the integral $\lambda_{{}_{{\widehat{(H,\mathcal{B})}}_{ad}}}$ of the adjoint hypergroup ${\widehat{(H,\mathcal{B})}}_{ad}$.
@@
For any weakly-integral premodular category $\mathcal{C}$, ${\widehat{{K(\mathcal{C})}}}$ is both RN and rational, as demonstrated in [Bur21, Theorem 1.2]. Furthermore, for any unitary fusion category $\mathcal{C}$, ${\widehat{{K(\mathcal{C})}}}$ is also RN, see [LPW21, ENO21], and primary $3$-criterion in [HLPW24]. A normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ is called dual-Burnside if its dual ${\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is Burnside (see examples from group theory in §12). By duality, one obtains the following:
@@
The following are equivalent for an abelian RN-hypergroup ${(H,\mathcal{B})}$.
@@
    (1.5) $\big{(}\prod_{i=1}^{m}{\frac{x_{i}}{d_{i}}}\big{)}^{2}=\frac{1}{n({{(H,\mathcal{B})}_{ad}})}\big{(}\sum_{x_{i}\in\mathcal{B}_{ad}}d_{i}x_{i}\big{)}$
@@
Let ${(H,\mathcal{B})}$ be a dualizable ARN-hypergroup. Then ${(H,\mathcal{B})}$ is nilpotent if and only if its dual ${\widehat{(H,\mathcal{B})}}$ is also nilpotent. Moreover, in this case they have the same nilpotency class.
@@
Note that the above theorem holds for weakly-integral modular categories since their Grothendieck rings are Burnside by [Bur23, Theorem 2].
@@
Concrete examples on which these results apply can be found in §12.3. Next theorem improves [OY23, Theorem 4.5].
@@
Let $\mathcal{C}$ be a weakly-integral modular fusion category of $\mathrm{FPdim}(\mathcal{C})=dm$ with $d$ square-free coprime with $m$ and $d_{i}^{2}$ for all $i$. Then $d\mid\mathrm{FPdim}({\mathcal{C}_{\mathrm{pt}}})$ and $\mathcal{C}$ admits a decomposition
@@
where ${\mathcal{D}}$ is a pointed non-degenerate braided fusion category of dimension $d$ and ${\mathcal{D}}^{\prime}$ a non-degenerate braided fusion category of dimension $m$.
@@
Let $\mathcal{C}$ be any integral modular fusion category of $\mathrm{FPdim}(\mathcal{C})=dm$ with $d$ square-free coprime with $m$. Then $d\mid\mathrm{FPdim}({\mathcal{C}_{\mathrm{pt}}})$ and $\mathcal{C}$ admits a decomposition
@@
where ${\mathcal{D}}$ is a pointed non-degenerate braided fusion category of dimension $d$ and ${\mathcal{D}}^{\prime}$ a non-degenerate braided fusion category of dimension $m$.
In particular, the $\mathrm{FPdim}$ of a perfect integral modular fusion category does not have a powerless prime factor (Corollary 10.11). In particular, a perfect modular fusion category of even $\mathrm{FPdim}$ has its $\mathrm{FPdim}$ divisible by $4$.
@@
By the classification of non-pointed integral modular fusion categories of small rank in [ABPP] and by Corollary 1.18, the $\mathrm{FPdim}$ of a non-pointed integral modular fusion category does not have a powerless prime factor for every rank less than $11\times 3=33$, and every rank less than $17\times 7=119$ in the odd-dimensional case.
Integral modular fusion categories were recently intensively studied in the literature. In [CP22] it was formulated the following conjecture
@@
There is no perfect modular fusion category of odd $\mathrm{FPdim}$.
Note that the above result also applies to non-pointed simple integral modular fusion categories (since they are in particular perfect). The importance of the existence of such categories was described in [LPR23, §5], where the following two open statements are proved to be equivalent, whereas Proposition 10.14 states that they imply Conjecture 1.20.
Statement 1.21.
@@
Shortly, the organization of this paper is the following. In §2 we recall the basics on hypergroups and abelian rational normalizable hypergroups. In §3 we develop few basic facts about Galois symmetries of rational hypergroups. In §4 hypergroups with Burnside property are studied and Theorem 1.7 is proven. In the same section we prove Theorem 1.2 and its more general version Theorem 4.17 for hypergroups. In this section, Burnside’s property for dual hypergroups is also studied and Theorem 1.8 is proven. In §5 we extend the notion of kernels of objects from the settings of fusion categories to the settings of ARN-hypergroups. In §7 we construct the universal grading group for ARN-hypergroups. A connection between this universal grading group and the group of grouplike elements of the dual it is proven in Theorem 6.5. §7 describes the constriction of the universal grading group of an ARN-hypergroup. In §8 we define nilpotent ARN-hypergroup and prove Theorems 1.11 and 1.12. The applications to modular fusion categories are given in §10.1. In §10 we prove Proposition 10.14 showing that any of the two statements from [LPR23, §5] implies Conjecture 1.20.
Next, §11 delves into applications and supplementary results: §11.1 offers a characterization for a Drinfeld center to be perfect; §11.2 establishes some results concerning integrality and Burnside properties, proving Theorems 1.10 and 1.23, particularly highlighting their equivalence in the perfect modular case; §11.3 characterizes the near-group fusion categories which admit a modular structure.
Finally, §12 presents concrete examples and counterexamples: §12.1 introduces non-Burnside simple integral fusion rings; §12.2 classifies certain dual-Burnside $\mathrm{Rep}(G)$; §12.3 outlines the classification of fusion rings lacking modular categorification within certain families, employing skills from this paper.

### v11

@@ Abstract.
In this paper, we extend a classical vanishing result of Burnside from the character tables of finite groups to the character tables of commutative fusion rings, or more generally to a certain class of abelian normalizable hypergroups. We also treat the dual vanishing result. We show that any nilpotent unitary fusion categories satisfy both Burnside’s property and its dual. Using Drinfeld’s map, we obtain that the Grothendieck ring of any weakly-integral modular fusion category satisfies both properties. As applications, we prove new identities that hold in the Grothendieck ring of any weakly-integral fusion category satisfying the dual-Burnside’s property, thus providing new categorification criteria. In particular we improve [OY23, Theorem 4.5] as follows: A weakly integral modular fusion category of $\mathrm{FPdim}$ $md$ with $d$ square-free coprime with $m$ and $\mathrm{FPdim}(X)^{2}$ for every simple object $X$, has a pointed modular fusion subcategory of $\mathrm{FPdim}$ $d$. We also present new results on perfect modular fusion categories, including a Cauchy-type theorem.
@@
18M20; 20N20; 20C15; 18N25; 16T20; 16T30
The first author is supported by a grant of the Ministry of Research, Innovation and Digitization, CNCS/CCCDI - UEFISCDI, project number PN-III-P4-ID-PCE-2020-0878, within PNCDI III. The second author is supported by BIMSA Start-up Research Fund, Foreign Youth Talent Program from the Ministry of Sciences and Technology of China and National Natural Science Foundation of China (NSFC, Grant no. 12471031)
@@ 1. Introduction
A classical result of Burnside in the character theory of finite groups states that any irreducible non-linear character of a finite group vanishes on at least one element of the group. This can be stated as follows: in the character table of a finite group, the row of every character of degree 1$\neq$ 1≠ 1 contains a zero entry.
More recently, the dual version of this result was also intensively studied in the literature. This dual version describes the group elements that vanish on at least one irreducible character, see [INW99] and the references therein. Burnside’s result was extended to every weakly integral fusion category with a commutative Grothendieck ring, initially in the context of modular categories in [GNN09, Appendix], and subsequently in the general framework in [Bur23].
@@
Recall that in [Bur21] the author introduced the notion of dual of a Grothendieck ring. This was achieved based on work of Harrison [Har79] on dualizable probability groups and on [ZZ21]. More generally, this notion of dual of a fusion ring was extended to arbitrary fusion rings in [BP25]. It was noticed in [Bur21] that the dual of the Grothendieck ring of a pivotal fusion category is isomorphic to the center of the category as defined in [Shi17].
@@
(1.1) $aa^{*}=N_{a,a^{*}}^{1}1\text{ (or equivalently, }h_{a}aa^{*}=1\text{, with }h_{a}:=1/N_{a,a^{*}}^{1})$
is called a grouplike element of ${(H,\mathcal{B})}$. In other words, $N_{a,a^{*}}^{c}=\delta_{1,c}N_{a,a^{*}}^{1}$.
@@ 1.1. The dual hypergroup of a normalized hypergroup
Let ${(H,\mathcal{B})}$ be an abelian normalized hypergroup. By the abelian assumption, $^{\widehat{\mathcal{B}}}$ forms a basis for the dual $H^{*}$, which then is $\mathbb{C}[{\widehat{\mathcal{B}}}]$, and on which we can define a multiplication. On the basis $^{\widehat{\mathcal{B}}}$, the multiplication $\mu_{i}\star\mu_{j}$ is defined by declaring
@@
and then extending linearly $\mu_{i}\star\mu_{j}$ on the entire ${\mathbb{C}[\mathcal{B}]}$. In this way, we obtain an algebra structure on $H^{*}$.
@@
By [Bur21, Proposition 2.3], for any abelian normalized hypergroup $(H,{\mathcal{B}})$, the pair $(H^{*},{\widehat{\mathcal{B}}})$ is also an abelian normalized hypergroup. Moreover, the involution on $^{\widehat{\mathcal{B}}}$ is given by $\mu_{j}\mapsto\mu_{j^{\#}}$, where $\mu_{j^{\#}}\in{\widehat{\mathcal{B}}}$ is defined by $\mu_{j^{\#}}(a):=\mu_{j}(a^{*})$ for all $a\in\mathcal{B}$. We denote by $\widehat{{(H,\mathcal{B})}}:=(H^{*},{\widehat{\mathcal{B}}})$ the dual hypergroup of the abelian normalized hypergroup ${(H,\mathcal{B})}$. It is also well known that for a normalized hypergroup ${(H,\mathcal{B})}$, then $\mu_{j}(a^{*})=\overline{\mu_{j}(a)}$ for all $a\in\mathcal{B}$. We also use the notation ${\widehat{h}}_{j}:={\widehat{p}}_{1}(j,j^{\#})^{-1}$, called the order of $\mu_{j}$ in the dual hypergroup $^{\widehat{(H,\mathcal{B})}}$, see Lemma 2.8.
For any RN-hypergroup ${(H,\mathcal{B})}$, we denote by ${(H,\mathcal{B})}_{ad}$ its adjoint sub-hypergroup, see §6.2. As for fusion rings, this is defined as the sub-hypergroup of ${(H,\mathcal{B})}$ generated by the element $I(1):=\sum_{i\in\mathcal{I}}h_{i}x_{i}x_{i^{*}}$, where $h_{i}:=1/N^{1}_{i,i^{*}}$.
@@
A hypergroup ${(H,\mathcal{B})}$ is called normalizable if we obtain a normalized hypergroup after rescaling it. (see §2.2 for more details.)
@@ 1.2. Main results of the paper
Let $(H,B,\mu_{1})$ be an abelian normalizable hypergroup such that $G{(H,\mathcal{B})}$ is a finite group. Then $(H,{\mathcal{B}})$ is Burnside if and only if
$(\prod_{j=1}^{m}\mu_{j})^{2}=\frac{|G{(H,\mathcal{B})}|}{n{(H,\mathcal{B},\mu_{1})}}\big{(}\sum_{j\in\mathcal{I}_{G{(H,\mathcal{B})}}}{\widehat{h}}_{j}\mu_{j}\big{)}$
See Definition 2.5 for $n(\_)$). We characterize Burnside dualizable ARN-hypergroups as follows:
Theorem 1.8.
Let ${(H,\mathcal{B},\mu_{1})}$ be an abelian normalizable hypergroup such that $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is RN. Then $(H,{\mathcal{B}})$ is Burnside if and only if the following identity holds on $^{\widehat{(H,\mathcal{B})}}$:
@@
Note that the RHS of the equation above corresponds to the integral $\lambda_{{}_{{\widehat{(H,\mathcal{B})}}_{ad}}}$ of the adjoint hypergroup ${\widehat{(H,\mathcal{B})}}_{ad}$.
Remark 1.9.
If the abelian normalizable hypergroup ${(H,\mathcal{B})}$ is either RN or dual RN, then according to Lemmas 2.32 and 2.36, the group $G{(H,\mathcal{B})}$ must be finite. As a result, we can apply Theorem 1.7. Specifically, by comparing this with Theorem 1.8, we can infer that in the dual RN case, $n(\widehat{(H,{\mathcal{B}})}_{ad})|G{(H,\mathcal{B})}|=n{(H,\mathcal{B})}$. By duality, we also have $n((H,{\mathcal{B}})_{ad})|G{\widehat{(H,\mathcal{B})}}|=n{(H,\mathcal{B})}$ in the RN case. These equalities can also be obtained from Harrison’s Equations (7.7) and (7.6), although in the more restrictive dualizable ARN case.
@@
For any weakly-integral premodular category $\mathcal{C}$, $^{\widehat{{K(\mathcal{C})}}}$ is both RN and rational, as demonstrated in [Bur21, Theorem 1.2]. Furthermore, for any unitary fusion category $\mathcal{C}$, $^{\widehat{{K(\mathcal{C})}}}$ is also RN, see [LPW21, ENO25], and primary 333-criterion in [HLPW24]. A normalizable hypergroup ${(H,\mathcal{B},\mu_{1})}$ is called dual-Burnside if its dual $^{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is Burnside (see examples from group theory in §12). The dual version of Theorem 1.7, is the following:
@@
Let $(H,B,\mu_{1})$ be an abelian normalizable hypergroup such that $G{\widehat{{(H,\mathcal{B},\mu_{1})}}}$ is a finite group. Then $(H,{\mathcal{B}})$ is dual-Burnside if and only if
$(\prod_{i=1}^{m}{\frac{x_{i}}{d_{i}}})^{2}=\frac{|G{\widehat{{(H,\mathcal{B},\mu_{1})}}}|}{n{(H,\mathcal{B},\mu_{1})}}\big{(}\sum_{i\in\mathcal{I}_{G{\widehat{{(H,\mathcal{B},\mu_{1})}}}}}h_{i}d_{i^{*}}x_{i}\big{)}$
On the other hand, the dual version of Theorem 1.8 is:
Theorem 1.11.
The following are equivalent for an ARN-hypergroup ${(H,\mathcal{B})}$.
@@
    (1.5) $\big{(}\prod_{i=1}^{m}{\frac{x_{i}}{d_{i}}}\big{)}^{2}=\frac{1}{n({{(H,\mathcal{B})}_{ad}})}\big{(}\sum_{x_{i}\in\mathcal{B}_{ad}}h_{i}d_{i^{*}}x_{i}\big{)}$
@@
Let ${(H,\mathcal{B})}$ be a dualizable ARN-hypergroup. Then ${(H,\mathcal{B})}$ is nilpotent if and only if its dual $^{\widehat{(H,\mathcal{B})}}$ is also nilpotent. Moreover, in this case they have the same nilpotency class.
@@
Note that the above theorem holds for weakly-integral modular categories since their Grothendieck rings are Burnside by [GNN09, Appendix] or [Bur23, Theorem 2].
@@
In particular,
Remark 1.20 (Cauchy-type theorem).
For every perfect integral modular fusion category,
(1.15) $\mathcal{V}(\mathcal{C})=\bigcup_{i=1}^{m}\mathcal{V}(d_{i}),$
thus, for all prime $p$ dividing $\mathrm{FPdim}(\mathcal{C})$, there is a simple object $X$ in $\mathcal{C}$ such that $p$ divides $\mathrm{FPdim}(X)$; which can be interpreted as a Cauchy-type theorem. Consequently, $p^{2}$ divides $\mathrm{FPdim}(\mathcal{C})$ since $\mathrm{FPdim}(X)^{2}$ divides $\mathrm{FPdim}(\mathcal{C})$ by [EGNO15, Proposition 8.14.6].
Concrete examples where these results apply can be found in §12.3. Next theorem improves [OY23, Theorem 4.5].
@@
Let $\mathcal{C}$ be a weakly-integral modular fusion category of $\mathrm{FPdim}=dm$ with $d$ square-free coprime with $m$ and $d_{i}^{2}$ for all $i$. Then $d\mid\mathrm{FPdim}({\mathcal{C}_{\mathrm{pt}}})$ and $\mathcal{C}$ admits a decomposition
@@
where ${\mathcal{D}}$ is a pointed modular fusion category of dimension $d$ and ${\mathcal{D}}^{\prime}$ a modular fusion category of dimension $m$.
@@
Let $\mathcal{C}$ be any integral modular fusion category of $\mathrm{FPdim}=dm$ with $d$ square-free coprime with $m$. Then $d\mid\mathrm{FPdim}({\mathcal{C}_{\mathrm{pt}}})$ and $\mathcal{C}$ admits a decomposition
@@
where ${\mathcal{D}}$ is a pointed modular fusion category of dimension $d$ and ${\mathcal{D}}^{\prime}$ a modular fusion category of dimension $m$.
In particular, the $\mathrm{FPdim}$ of a perfect integral modular fusion category does not have a powerless prime factor (Corollary 10.10). In particular, a perfect modular fusion category of even $\mathrm{FPdim}$ has its $\mathrm{FPdim}$ divisible by 444.
@@
By the classification of non-pointed integral modular fusion categories of small rank in [ABPP] and by Corollary 1.22, the $\mathrm{FPdim}$ of a non-pointed integral modular fusion category does not have a powerless prime factor for every rank less than 11×3=3311$\times$ 3=3311 × 3 = 33, and every rank less than 17×7=11917$\times$ 7=11917 × 7 = 119 in the odd-dimensional case.
Integral modular fusion categories were recently intensively studied in the literature. The following conjecture is formulated in [CP22]:
@@
There is no non-trivial perfect modular fusion category of odd $\mathrm{FPdim}$.
This conjecture was checked for ranks less than 252525 in [ABPP] or [CGP23]. Note that the above result also applies to non-pointed simple integral modular fusion categories (since they are in particular perfect). The importance of the existence of such categories was described in [LPR23, §5], where the following two open statements are proved to be equivalent, whereas Proposition 10.13 states that they imply Conjecture 1.24.
Statement 1.25.
@@
Shortly, the organization of this paper is the following. In §2 we recall the basics on hypergroups and abelian rational normalizable hypergroups. In §3 we develop few basic facts about Galois symmetries of rational hypergroups. In §4 hypergroups with Burnside property are studied. In the same section, we prove Theorem 1.2 and its more general version Theorem 4.17 for hypergroups. In this section, Burnside’s property for dual hypergroups is also studied. In §5 we extend the notion of kernels of objects from the settings of fusion categories to the settings of ARN-hypergroups. In §6 and §7, we construct the universal grading group for ARN-hypergroups. A connection between this universal grading group and the group of grouplike elements of the dual is proven in Theorem 6.5. Theorems 1.8 and 1.11 are proven. In §8 we define nilpotent ARN-hypergroup and prove Theorems 1.14 and 1.15. In §9 we prove Theorem 1.12. In §10, the applications to modular fusion categories are given (e.g. the proofs from Theorem 1.16 to Corollary 1.22), and we prove Proposition 10.13 showing that any of the two statements from [LPR23, §5] implies Conjecture 1.24.
Next, §11 delves into applications and supplementary results: §11.1 offers a characterization for a Drinfeld center to be perfect; §11.2 establishes some results concerning integrality and Burnside properties, proving Theorems 1.13 and 1.27, particularly highlighting their equivalence in the perfect modular case; §11.3 characterizes the near-group fusion categories which admit a modular structure.
Finally, §12 presents concrete examples and counterexamples: §12.1 introduces non-Burnside simple integral fusion rings; §12.2 classifies certain dual-Burnside $\mathrm{Rep}(G)$; §12.3 provides certain families of fusion rings lacking modular categorification by Theorems 1.19 or 1.22.